OpenStudy (gio333):

What is the restriction on the rule that says, when two numbers are multiplied within a square root, you can split it into a multiplication of two square roots?

OpenStudy (ghazi):

you've to read properties of surds

OpenStudy (anonymous):

\[\sqrt{xy} =\sqrt{x}\sqrt{y}\] \[x \ge 0 and y \ge0\]